OpenStudy (anonymous):

if pq=12 and (p+q)2=24, then find the value of p2+q2. The 2 are the square units

OpenStudy (anonymous):

Use ^ for the power (p+q)^2 = p^2 + 2pq + q^2 = 24 p^2 + q^2 = 24 - 2pq pq = 12 Therefore p^2 + q^2 = 24 -24 = 0